OpenStudy (dmezzullo):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = −14 x = −17

OpenStudy (dmezzullo):

@sue101 @rduncan16 @mikala1 @hero @Ledah @SUpAmArIo @jazy @jhonyy9

OpenStudy (anonymous):

ur so needy danny :P

OpenStudy (anonymous):

(x + 20)/2 = 3x x + 20 = 3x * 2 = 6x x - 6x = -20 -5x = -20 x = -20/-5 = 4 x = 4

OpenStudy (anonymous):

\[\frac{ x + 20 }{ 2 } = 3x\]Multiply both sides by 2.\[\frac{ x + 20 }{ 2 }(2) = 3x(2)\]\[x + 20 = 6x\]Combine like terms and solve for x.

jhonyy9 (jhonyy9):

cross multiplie

OpenStudy (anonymous):

x + 20/2 = 3x multiple both sides by 2 to get rid of the 2 2(x+20/2) = 3x(2) the 2s cancel each other out on the left side x+20=6x you want to get the x's on the same side so subtract x from each side 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

jhonyy9 (jhonyy9):

what will get than ?

OpenStudy (dmezzullo):

cse im so stupid @sue101

OpenStudy (anonymous):

opp i put the answer on accsident

OpenStudy (dmezzullo):

4x

OpenStudy (dmezzullo):

cse 20/5=4

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, x = 4.